5. A steady state, steady flow compressor draws in 240 liters per second of air whose density is 1.26

kg m /m 3 and discharges it with a density of 4.9 kg m /m 3 . At suction P 1 = 104 kPaa; at discharge, P 2 = 551

kPaa. The increase of specific internal energy is 78.5 kJ/kg m and the heat from the air by cooling is 30

kJ/kg m . Neglecting the change in potential and kinetic energies, determine the work in kJ/min.


1
Expert's answer
2021-09-16T18:17:24-0400

"U=um_1=u\rho_1V,"

"Q=qm_2=q\rho_2V,"

"a=\frac A{60}=\frac{U+Q}{60}=\frac{(u\rho _1+q\rho _2)V}{60}=984~\frac{kJ}{min}."